width of a triangle is 2ft less than the length, area is 24f^2, L=, w=
Answer by checkley77(12844)

W=L-2
LW=AREA
L(L-2)=24
L^2-2L-24=0
(L-6)(L+4)
L-6=0
L=6 ANS. FOR THE LENGTH.
W=6-2=4 ANS. FOR THE LENGTH.
PROOF:
6*4=24
24=24